# ”Living Within Four Walls”: Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement

### v2

@@ 3.1. Background
To prevent the community spread of the COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (58; 85). Our data collection was conducted during the second lockdown period with Stage IV restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1). Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage IV restrictions in metropolitan Melbourne (62).
Under the Stage IV restrictions in force at the time of our data collection, a daily curfew was implemented from 8 PM to 5 AM, forbidding anyone from leaving their homes except for special reasons. Residents were only allowed to shop and exercise within 5 km of their homes, and all students returned to home-based learning. Other restrictions that were previously applied only to specific districts were extended to the entire Melbourne metropolitan area. On 13 September 2020, several restrictions were eased in Melbourne, including reduced curfews and loosened rules on outdoor exercise and social interactions. During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (74; 82). This presents a unique opportunity for researchers to gain insights into the psychological states and smartphone usage behaviours of individuals during home confinement. The knowledge gained could prove invaluable in understanding individuals in similar home confinement situations, such as those who are ill or undergoing post-recovery home rehabilitation, and home-based work.
(a) Age and gender
(b) Smartphone and desktop
Figure 2. Demographic and device usage information of participants
3.2. Participants
Participants were recruited through advertising and comprised of university students, professional workers, academic staff, and health workers. In total, we recruited 32 participants (17 males and 15 females), where most participants were between 18-40 years old (mean = 25.44, median = 23.00, std = 8.95). The age distribution of participants across genders can be found in Figure 2. It is worth noting that participants were primarily recruited through advertisements on social media platforms Facebook and Discord, which may have resulted in an over-representation of users of these platforms in our sample. While this could be considered a limitation of our study, it is important to note that our research did not specifically focus on social media usage behaviour. Rather, we focus on individuals’ psychological states and smartphone usage behaviours during COVID-19 home confinement. Therefore, we believe that the potential oversample threat to the validity of our results is minimized. However, it is still important to note that our findings are specific to this particular population subset and may not be necessarily generalizable to other populations. Future research should consider recruiting participants from a wider range of sources to improve the generalizability of the findings.
After the data collection, each participant was compensated with a 25 AUD gift card as a token of appreciation for their time. Of the 32 participants, 27 recorded smartphone usage, 13 recorded desktop usage, and 8 recorded both smartphone and desktop app usage. Due to the relatively small number of participants using desktops, we primarily focused on analyzing the smartphone usage of the 27 participants. Analysis of desktop usage behaviours will be explored further in future work.
@@ 3.3. Procedures
Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop applications (36) on their smartphones and desktops, respectively. These applications were specifically designed to record usage behaviours and enable background sensing and experience sampling, drawing inspiration from Van et al. (90). Both applications were optimized to minimize their consumption of computational power, energy, and storage resources. The applications utilized ESM, popping up a short questionnaire to participants every 90 minutes when their devices were actively in use between 7 am and 10 pm. This helps to obtain unbiased samples of participants’ activities, which served as the ground truth for our predictive models. Furthermore, an event-based ESM was triggered on the user’s smartphone after a minimum interaction of 10 minutes, with subsequent event-based ESMs occurring every 30 minutes.
At the end of each day, participants received an end-of-day survey via email at 7:30 pm. This survey aimed to gather reflections on their well-being and the impact of the lockdown on their daily experiences. This provides valuable context and additional data beyond the ESM questionnaires. Overall, we collected 502,485 recordings for smartphone usage, 1,749 ESM responses, and 265 end-of-day responses.
Table 1. An overview of the collected data
@@
Smartphone Usage Application usage Foreground application name, package identifier, category N/A
@@
3.3.1. Smartphone usage data
During the data collection, the background services of smartphones and desktops continuously tracked the following information: application usage (using Android Accessibility Services  [^2] Accessibility Services: https://developer.android.com/reference/android/accessibilityservice/AccessibilityService), notifications (using Android Notification Listeners  [^3] Notification Listeners: https://developer.android.com/reference/android/service/notification/NotificationListenerService), as well as the phone’s states (i.e. screen/battery status, ringer modes, location updates and physical activities). Specifically, the mobile app usage behaviour was recorded every time the user opened the app. For each notification, the application recorded the arrival time, hashed content (used for calculating the length of notification), the sender (private, work only, both or none), and participants’ relationship to the sender (family, friend, work and none). Table 1 displays an overview of the collected data for smartphone usage, ESM and end-of-day survey. In total, we have collected 502,485 recordings for smartphone usage data from 27 participants.
3.3.2. ESM responses
There are some popular questionnaires to measure well-being (43; 10), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (10) was employed in this research. In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and task from users. Especially, the item ‘valence’ was rated with a 5-point Likert scale from ‘unhappy’ to ‘happy’, and the item ‘arousal’ was also rated with a 5-point Likert scale from ‘calm’ to ‘excited’. Both valence and arousal were used to represent the users’ emotions based on the circumplex model (79). Understanding the social role, interruptiblity and task help us understand how people manage their time and attention on the digital devices (3). The detailed items of the ESM can be found in Table 1 and the screenshot of the ESM on users’ smartphones was shown in Figure 3(a). Overall, we have collected 1,749 ESM responses from 29 participants. Of these, 1,522 responses were obtained from 25 participants using smartphones (2 participants did not contribute any ESM responses), while the remaining 227 responses were from 10 participants using desktops.
@@ 3.3.3. End-of-day survey
As shown in Figure 3(b), the end-of-day surveys were optional and included ten items to understand participants’ well-being during the lockdown such as productivity, valence, energy, and work-life balance. In addition, we surveyed how participants feel today compared to yesterday. This information could describe the dynamics of people’s experience during the COVID-19 lockdown. In the end-of-day survey, each item was rated with a 5-Likert scale or answered with free text input (see Table  1). Specifically, item 1 and item 2 were rated from 1 to 5 which indicates ‘strongly agree’, ‘somewhat agree’, ‘neither agree nor disagree’, ‘somewhat disagree’, ‘strongly disagree’, and items 3, 4, 5, 9 were rated from 1 to 5 which indicates ‘a great deal’, ‘a lot’, ‘a moderate amount’, ‘a little’, ‘not at all’. In total, we collected 265 responses from 25 participants during the data collection period.
@@
4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement
In this section, we aim to answer RQ1 ”What are people’s emotions and social roles like during COVID-19 home confinement? To what extent does the COVID-19 lockdown influence people’s emotions and work-life balance in their daily lives?”. To achieve this, we analyse 1522 ESM responses collected through smartphones from 25 participants, as described in Section 4.1. Furthermore, we explore the influence of the COVID-19 lockdown by examining the end-of-day survey in Section 4.2.
@@ 4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement
4.1. The Perceived Valence, Arousal and Social roles
The distribution of perceived valence, arousal, and social roles from the ESM responses is depicted in Figure 4. On average, participants reported a valence level of 3.46 (SD = 1.01) and an arousal level of 2.72 (SD = 1.11). Usually, the relatively high valance and low arousal refer to a state of calmness, relaxation and peace. This could be attributed to various factors during the COVID-19 home confinement. One possible reason is that people may feel relieved to be safe at home, and cherish the opportunity to spend more quality time with their family and loved ones, away from the potential risks of infection outside. Furthermore, the absence of the usual stressors of daily life, such as commuting, work pressures, and social obligations, may be contributing to a general sense of calm and well-being. Additionally, some people may have found comfort in the sense of community and solidarity that emerged during the pandemic, as people came together to support one another through difficult times. All of the factors mentioned earlier may contribute to the high valence and low arousal state during COVID-19 home confinement. Nevertheless, it is crucial to acknowledge that our data collection was limited to a three-week period, as COVID-19 Stage IV restrictions (strict home confinement) were lifted and steps for reopening commenced. Thereafter, in the event that strict home confinement lasts longer, it is possible that the emotional experiences of individuals may evolve as they adapt to the new normal and encounter new challenges. For instance, prolonged periods of confinement may lead to feelings of boredom, loneliness, and frustration.
@@ 4.1. The Perceived Valence, Arousal and Social roles
Next, we investigate the distribution of social roles and their correlation with emotion during home confinement. A social role refers to a set of norms, expectations, rules, beliefs and behaviours that are associated with a particular social situation (75). These roles can significantly influence users’ behaviours and the type of information they prefer to receive (3). As shown in Figure 4(b), smartphone users predominantly engaged in a private-related role during home confinement, which suggests that individuals tend to prioritize their personal lives and relationships while staying at home.
Since social roles play a crucial role in shaping individuals’ well-being, we investigate the emotions of users in different social roles, as shown in Figure 4(c). We found that, when engaging in private-related roles, participants had the highest levels of valence and lowest levels of arousal compared to work-related roles and hybrid roles. This may be because private-related roles involve personal activities which can provide a sense of comfort and happiness, while work-related roles or hybrid roles may involve stress, deadlines, and pressure to perform. The results of Welch’s t-tests (99) further support these findings, indicating that there is a significant difference in valence levels between private-related roles and the other two roles (i.e., private vs. both, private vs. work), with p-value less than 0.05. The above findings provide insights into how individuals can structure their daily routines to promote positive emotions during periods of home confinement, and interventions aimed at improving well-being during home confinement may benefit from considering the types of social roles that individuals are engaging in (34).
4.2. The Influence of the COVID-19 Lockdown on Emotions, Social Roles and Work-life Balance
In total, 54.3%, 51.3% and 54.7% of end-of-day responses displayed that the lockdown has influenced their valence, energy, and work-life balance to some extent, as shown in Figure 5. For the influence of lockdown on valence, 13 participants reported that it had a negative effect, claiming they ‘felt quite anxious and sad’ as ‘staying at home has made me a little sad’. They complained that they were unable to see friends (9 responses), see families (5 responses), go outside (10 responses), and feel bored (6 responses), e.g., ‘I thought I might be able to see friends, but I did not’, ‘I want to go outside, to the park’, ‘I’m missing my family’, ‘It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long. I’m bored and sick of my housemates and it feels like time has stopped ’. Nevertheless, two felt the lockdown had a positive effect, claiming that ‘I am a lot happier’ and ‘I was able to maximize my time to do some exercise and transition to work immediately’.
@@ 4.2. The Influence of the COVID-19 Lockdown on Emotions, Social Roles and Work-life Balance
Regarding the impact of lockdown on arousal, 11 participants claimed that the impact was negative, reporting feeling tired (11 responses), being lazy (5 responses) and losing motivation (3 responses), e.g., ‘I don’t feel as motivated to do productive things’, ‘Emotionally tired, perhaps as a result of the lockdown’. However, two participants claimed the lockdown gave them ‘more energy’. For the influence on work-life-balance, most participants reported that they faced more interruptions or distractions (14 responses), had difficulty separating work and life (11 responses), spent more time on work due to interruptions or blurring of personal and work boundaries (15 responses), e.g., ‘I worked a bit more than I should’ve’, ‘Everything blends together, separating work and life is really difficult’, ‘Continuous interruptions, even my cat interrupts me’, ‘Need to work longer hours due to interruptions’. Overall, while a minority of responses showed that lockdown has a positive impact on their lives (e.g., more time with family or gardening), most responses showed negative influences of lockdown on their valence, arousal and work-life balance.
5. Understanding App Usage Behaviours During COVID-19 Home confinement
In this section, we address RQ2 ”Is it possible to unsupervisedly extract user activity from app usage records? If so, how are these activities related to emotions, productivity and interruptions?”. First, we introduce the process of app categorization. Next, we conduct unsupervised activity recognition from app sequences and display the average activities for all clusters. In this end, we explore the correlation between user activities and perceived psychological metrics.
Table 2. App categories and top three frequently used apps
Upper Category Category Top App Second App Third App
@@ 5.1. App Categorization
We categorized the apps using historical usage logs. For Android apps, we referred to the categories listed in the Google Play Store, excluding uncategorized apps. Windows apps were manually sorted into the same categories as Google Play Store apps. Less frequently used apps, with a cumulative usage time of less than one hour across all participants, were excluded. Overall, we recorded 28 app categories and consolidated them into 8 upper-level categories: Business, Communication, Entertainment, Productivity, Reading, Social, Tools and Others. A summary of the top three most frequently used apps in each category is provided in Table 2. Notably, Chrome is categorized as a Communication app due to the scarcity of browser apps in the Google Play Store.
Categorizing desktop apps within the same categories used for mobile apps proves challenging, as they are often used for work and software development (e.g., Pycharm, WinSCP etc). In light of this, we grouped several development and utility apps under the Tools Category. Overall, the top 10 Tools apps for both desktop and mobile include: Gboard, Explorer, Samsung Keyboard, Always on Display, WinSCP, Pixel Launcher, Acrobat, pychar64, Google and Samsung keyboard. It is worth mentioning that developer tools such as Pycharm and WinSCP, as well as the utility Explorer, fall under this category.
5.2. Unsupervised Activity Recognition From App Sequences

### v3

@@ 3.1. A Case Study of COVID-19 Lockdown in Melbourne, Australia
To prevent the community spread of COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (47; 73). Our data collection was conducted during the second lockdown period with Stage IV restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1). Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage IV restrictions in metropolitan Melbourne (51).
Under the Stage IV restrictions in force at the time of our data collection, a daily curfew was implemented from 8 PM to 5 AM, forbidding anyone from leaving their homes except for special reasons. Residents were only allowed to shop and exercise within 5 km of their homes, and all students returned to home-based learning. Other restrictions that were previously applied only to specific districts were extended to the entire Melbourne metropolitan area. On 13 September 2020, several restrictions were eased in Melbourne, including reduced curfews and loosened rules on outdoor exercise and social interactions. During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (62; 70).
@@ 3.2. Procedures
Participants were recruited through advertising and comprised of university students, professional workers, academic staff, and health workers. In total, we recruited 32 participants (17 males and 15 females), where most participants were between 18-40 years old (mean = 25.44, median = 23.00, SD = 8.95). Of these, 27 participants provided smartphone data, while 13 contributed desktop data. Recruitment was primarily through social media platforms Facebook and Discord, potentially leading to an over-representation of users from these platforms. However, as our study focused on psychological states and digital device usage during COVID-19 home confinement, not social media behaviour, we believe this does not significantly impact the validity of our findings.
Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop apps(27; 80) on their smartphones and desktops. These apps, designed to record usage behaviors and enable experience sampling, were optimized for minimal computational, energy, and storage consumption. The apps used Experience Sampling Method (ESM) to prompt participants with a short questionnaire every 90 minutes during active device use between 7 am and 10 pm. To minimize participant burden, surveys were only triggered during active use. Additionally, the event-based ESM was triggered after 10 minutes of interaction, with follow-ups every 30 minutes. At day’s end, participants received an end-of-day survey via email at 7:30 pm. After the data collection, each participant was compensated with a 25 AUD gift card as a token of appreciation for their time.
Privacy, ethics, and safety. Our field study was designed and conducted after carefully reviewing potential risks and benefits, in compliance with the city’s COVID-19 safety regulations. The study has been approved by the Human Research Ethics Committee at the authors’ university. To ensure privacy, all personal identifying information was anonymized before data analysis. Participants’ consent was obtained prior to data collection, with a clear explanation of the study’s purpose and the nature of the data being collected. We also implemented strong data security measures, including encryption of sensitive data, to further protect participants’ privacy. Participants were informed of their right to withdraw at any time without penalty, and their data would be excluded from the study upon request.
Table 1. Overview of the collected data records across devices
Category Device Total Records Participants Data Used for RQs
App Both 502,485 32 RQ1
Desktop 261,911 13 /
Smartphone 240,574 27 RQ2
ESM Both 1,749 28 RQ1
Desktop 227 10 /
Smartphone 1,522 25 RQ2
Ene-of-Day Survey Both 265 25 RQ1
Table 2. Detailed overview of data categories, items, and response options
@@
Device Usage Application usage Foreground application name, package identifier, category N/A
@@
3.3. Collected Data
The data collection period spanned 21 days. From the 32 recruited participants, we collected 502,485 recordings of smartphone/desktop usage, 1,749 ESM responses, and 265 end-of-day surveys. Table 1 provides an overview of the collected data across devices. To address RQ1, we used both desktop and smartphone data to explore how well-being was during COVID-19 home confinement. For RQ2, due to the limited number of participants contributing desktop data, we focused exclusively on smartphone data and its corresponding ESM responses.
Smartphone/desktop usage data. During the data collection, background services of smartphones and desktops continuously tracked the application usage (via Android Accessibility Services  [^2] Accessibility Services: https://developer.android.com/reference/android/accessibilityservice/AccessibilityService), notifications (via Android Notification Listeners  [^3] Notification Listeners: https://developer.android.com/reference/android/service/notification/NotificationListenerService), as well as the phone’s states (only for smartphone, i.e. screen/battery status, ringer modes, location updates and physical activities). Specifically, smartphone app usage was recorded each time an app was opened, while desktop usage was tracked by polling the active application every second. For notifications, data captured included arrival time, hashed content (for notification length), sender (private, work, both, or none), and the participant’s relationship to the sender (family, friend, work, or none). Table 2 displays an overview of the collected data for digital device usage, ESM and end-of-day survey.
ESM responses. There are some popular questionnaires to measure well-being (34; 9), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (9) was employed in this research. In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and tasks from users (see Table 2). Specifically, the measurement of items valence and arousal were adapted from SAM  [^4] Compared to the original SAM questionnaire, we added a time constraint, asking participants to describe their feelings over the past hour. and rated with a 5-point Likert scale, from unhappy to happy for valence and from calm to excited for arousal. These measures were used to represent the users’ emotions based on the circumplex model (67). The social role, interruptibility, and task questions were sourced from Anderson’s assessment framework (4).
@@ 3.3. Collected Data
End-of-day survey. As shown in Figure 2(b), the end-of-day surveys were optional and included ten items designed to understand participants’ well-being during the lockdown, focusing on aspects such as productivity, valence, energy, and work-life balance. Additionally, we surveyed participants on how they felt today compared to yesterday. This information helps describe the dynamics of people’s experiences during the COVID-19 lockdown. The end-of-day survey questions were tailored to capture the unique context of our study and were not sourced from existing questionnaires. Each item in the end-of-day survey was rated using a 5-point Likert scale or answered with free text input (see Table  2). Specifically, items 1 and 2 were rated from 1 to 5, indicating strongly agree, somewhat agree, neither agree nor disagree, somewhat disagree, and strongly disagree. Items 3, 4, 5, and 9 were rated from 1 to 5, indicating a great deal, a lot, a moderate amount, a little, and not at all.
@@
4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement (RQ1)
In this section, we address RQ1 ”How are emotions and social roles during COVID-19 home confinement perceived through self-reports and inferred from app usage patterns?”. First, we analyse 1522 ESM responses collected via smartphones from 25 participants. Next, we explore the impact of the COVID-19 lockdown by examining the end-of-day survey. Additionally, we explore users’ app usage behaviours through unsupervised activity recognition, identifying common activity clusters and their relationship to the perceived psychological metrics.
@@ 4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement (RQ1)
4.1. The Perceived Valence, Arousal and Social Roles During Covid-19 Home Confinement
Figure 3 illustrates the distribution of perceived valence, arousal, and social roles. On average, participants reported a valence of 3.46 (SD = 1.01) and arousal of 2.72 (SD = 1.11), indicating a generally calm and relaxed emotional state. This likely reflects feelings of safety and relief from the external stressors associated with the pandemic, such as the risks of infection. The absence of daily stressors—like commuting, work pressure, and social obligations—may also contribute to this sense of well-being. Furthermore, the solidarity and support networks that emerged during the pandemic likely provided additional emotional comfort. However, it is important to note that our data collection occurred during a three-week period as the COVID-19 Stage IV restrictions were lifting.Had the confinement continued longer, participants likely would have reported more negative emotions, such as boredom or frustration, as they adjusted to the new normal.
Next, we investigate the role of social roles in shaping emotional experiences during home confinement. A social role refers to the expectations, norms, and behaviours associated with particular social contexts (63). These roles usually influence users’ behaviours and the type of information they prefer to receive (4). As shown in Figure 3(b), participants predominantly engaged in private-related roles, suggesting individuals tend to prioritize their personal lives and relationships during home confinement.
We further explore how emotions vary across different social roles. Figure 3(c) shows that participants reported the highest valence and lowest arousal when engaged in private-related roles, compared to work-related or hybrid roles. This suggests that personal activities, which are often less stressful and more enjoyable, foster a greater sense of well-being. Conversely, work-related and hybrid roles were associated with higher levels of stress and pressure. Welch’s t-tests (89) confirm these findings, showing significant differences in valence between private-related and the other two roles (i.e., private vs. both, private vs. work), with p-value ¡ 0.05. The findings offer insights into how individuals can structure their daily routines to boost positive emotions during home confinement. Interventions aimed at improving well-being in such situations may benefit from considering the social roles individuals are engaging in (26).
@@ 4.1. The Perceived Valence, Arousal and Social Roles During Covid-19 Home Confinement
4.2. The Influence of the Home Confinement on Emotions, Social Roles and Work-Life Balance
End-of-day survey responses reveal that the lockdown had a noticeable impact on participants’ valence, arousal, and work-life balance. As shown in Figure 4, 54.3%, 51.3%, and 54.7% of responses indicated that the lockdown influenced valence, energy, and work-life balance, respectively.
In terms of valence, 13 participants reported negative effects, describing feelings of anxiety, sadness, and boredom. Common complaints included missing social interactions (9 responses), the inability to go outside (10 responses), and a sense of boredom (6 responses). For instance, ”I thought I might be able to see friends, but I did not”, ”I want to go outside, to the park”, ”I’m missing my family”, ”It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long. I’m bored and sick of my housemates and it feels like time has stopped”. On the other hand, two participants reported positive effects, such as feeling happier or using the extra time for physical activity and seamless work transitions.
@@ 4.2. The Influence of the Home Confinement on Emotions, Social Roles and Work-Life Balance
Regarding arousal, 11 participants indicated a negative impact, expressing feelings of fatigue (11 responses), lack of motivation (3 responses), or laziness (5 responses), e.g., ”I don’t feel motivated to do productive things”, ”Emotionally tired, perhaps as a result of the lockdown”. However, two participants noted that they felt more energized during the lockdown.
In terms of work-life balance, the majority of participants reported they faced increased interruptions or distractions (14 responses), they had difficulty separating work from personal life (11 responses), and they spent more time working due to these disruptions (15 responses), with comments such as ”I worked a bit more than I should’ve”, ”Everything blends together, separating work and life is really difficult”. These findings suggest that the lockdown primarily had a negative impact on participants’ work-life balance, with many struggling to manage personal and professional responsibilities. While some respondents reported positive effects, such as spending more time with family or gardening, the overall trend indicates that the lockdown negatively influenced emotional well-being and work-life balance.
4.3. Understanding App Usage Behaviours During COVID-19 Home Confinement
Table 3. App categories (referred to the Google Play Store categories) and top three frequently used apps.
Upper Category Subcategory Top App Second App Third App
@@ 4.3.1. App Categorization
Apps were categorized based on historical usage logs. For Android apps, we used the categories from the Google Play Store, excluding uncategorized apps. Desktop applications were manually assigned to equivalent categories from the Google Play Store. Apps with less than one hour of cumulative usage across all participants were excluded. In total, we identified 28 app categories, which were consolidated into 8 high-level categories. A summary of the top three most frequently used apps in each category is provided in Table 3. Notably, Chrome was categorized under Communication, as the Google Play Store has few dedicated browser apps. Categorizing desktop applications posed a challenge, as many are used for work or software development (e.g., PyCharm, WinSCP). These were grouped under the Tools category, along with utility apps like Explorer.
4.3.2. Unsupervised Activity Recognition From App Sequences
To identify distinct user activities from app usage, we propose an unsupervised approach that combines segmentation and clustering. First, we segment the app usage sequences to capture shifts in activity, then cluster these segments to identify common patterns of behaviour.